# How Well Do LLMs Represent Values Across Cultures? Empirical Analysis of LLM Responses Based on Hofstede Cultural Dimensions

### v1

@@ Appendix B Hofstede Cultural Dimensions
There have been many attempts to define values that different cultures have. Going back to 1951, U.S. sociologists Talcott Parsons and Edward Shills defined cultural values as boiling down to choices between pairs of alternatives, including affectivity, self-orientation vs. collectivity-orientation, universalism, ascription, and specificity Parsons and Shils (1951). After greater improvements in the field of value collection from Florence Kluckhohn and Fred Strodtbeck Kluckhohn and Strodtbeck (1961), Mary Douglas Douglas (1973), Inkeles and Levinson Inkeles and Levinson (1969), Geert Hofstede Hofstede (1980) developed five unique cultural dimensions that take into account prior research on political systems (Gregg and Banks’ Gregg and Banks (1965)), economic development (Adelman and Morris’ Adelman and Morris (1967)), mental health (Lynn and Hampson’s Lynn and Hampson (1975)). Hofstede cultural dimensions are a way of defining values of different cultures based on pattern variables, or choices between pairs of alternatives. Although the data was initially collected in the 1980s, the validity of the cultural dimensions has held up to time as new data gets added (Hofstede and Bond (1988); Minkov (2007); Hofstede et al. (2010)). The most recent follow up studies have been in 2021 Minkov and Kaasa (2021), and 2022 Minkov and Kaasa (2022), showing that Hofstede cultural dimensions are relevant to the current day.
When considering other values to consider when analyzing LLMs, we examined GLOBE values – a large-scale study of leadership ideals, trust, and other cultural practices within 150 different countries – which build off the work of Hofstede cultural dimensions GLOBE Project (n.d.). However, while both Hofstede cultural dimensions, and GLOBE values have their origin in conducting research in the workforce, we found that GLOBE values are overly reliant on workforce and coworker/manager relations, and would not generalize as well to other, more diverse situations that values, such as Individualism vs. Collectivism could fall in. Furthermore, GLOBE values were supplied in ranges that are not as intuitive to understand, whereas Hofstede cultural dimensions are given as granular values, making it easier to compare values between countries.
@@ Appendix C Comparison Between Japanese and American Values
According to Hofstede cultural dimensions, Japan has an Individualistic vs. Collectivist score of 62, meaning that Japan is an individualistic country; in terms of granularity, Japan is more individualistic than the United States, which has an Individualistic vs. Collectivist score of 60. However, each LLM we tested along with each approach we tested perceived the United States as predominantly individualistic and Japan as predominantly collectivist, with the largest discrepancy being within the personas approach for Command-R, where 72.40% of responses for the American persona were individualistic and only 19.60% of answers for the Japanese persona where individualistic. This may be because much of English language data represents Japan as a collectivist country Scroope (2021) and the United States as an individualistic country Rosenbaum et al. (2018), leading to stereotypical representations of each country rather than true representations according to their Hofstede cultural dimensions. These findings hold for other individualistic countries often perceived as collectivist, such as South Korea Yuh (2016).
@@ Appendix E Full Data and Visualizations
Figure 3: GPT4o adhering well to individualism vs. collectivist value for high-resource languages
@@
Table 4: Language and Hofstede Cultural Dimensions Metadata
@@
Table 5: Graphs showing value differentiation across all models, approaches, and values. Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values.
@@
Table 6: Graphs showing value differentiation across all models, approaches, and values (continuation). Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
@@
Table 7: Graphs showing value differentiation across all models, approaches, and values (continuation). Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
@@
Table 8: Graphs showing value differentiation across all models, approaches, and values (continuation). Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
@@
Table 9: Graphs showing value differentiation across all models, approaches, and values. Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
@@
Table 10: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs.
@@
Table 11: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
@@
Table 12: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
@@
Table 13: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
@@
Table 14: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
@@
Table 15: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs.

### v2

@@ Appendix B Hofstede Cultural Dimensions
There have been many attempts to define values that different cultures have. Going back to 1951, U.S. sociologists Talcott Parsons and Edward Shills defined cultural values as boiling down to choices between pairs of alternatives, including affectivity, self-orientation vs. collectivity-orientation, universalism, ascription, and specificity (Parsons and Shils, 1951). After greater improvements in the field of value collection from Florence Kluckhohn and Fred Strodtbeck (Kluckhohn and Strodtbeck, 1961), Mary Douglas (Douglas, 1973), Inkeles and Levinson (Inkeles and Levinson, 1969), Geert Hofstede (Hofstede, 1980) developed five unique cultural dimensions that take into account prior research on political systems (Gregg and Banks’ (Gregg and Banks, 1965)), economic development (Adelman and Morris’ (Adelman and Morris, 1967)), mental health (Lynn and Hampson’s (Lynn and Hampson, 1975)). Hofstede cultural dimensions are a way of defining values of different cultures based on pattern variables, or choices between pairs of alternatives. Although the data was initially collected in the 1980s, the validity of the cultural dimensions has held up to time as new data gets added ((Hofstede and Bond, 1988); (Minkov, 2007); (Hofstede et al., 2010)). The most recent follow up studies have been in 2021 (Minkov and Kaasa, 2021), and 2022 (Minkov and Kaasa, 2022), showing that Hofstede cultural dimensions are relevant to the current day.
When considering other values to consider when analyzing LLMs, we examined GLOBE values – a large-scale study of leadership ideals, trust, and other cultural practices within 150 different countries – which build off the work of Hofstede cultural dimensions (GLOBE Project, nd). However, while both Hofstede cultural dimensions, and GLOBE values have their origin in conducting research in the workforce, we found that GLOBE values are overly reliant on workforce and coworker/manager relations, and would not generalize as well to other, more diverse situations that values, such as Individualism vs. Collectivism could fall in. Furthermore, GLOBE values were supplied in ranges that are not as intuitive to understand, whereas Hofstede cultural dimensions are given as granular values, making it easier to compare values between countries.
@@ Appendix C Comparison Between Japanese and American Values
According to Hofstede cultural dimensions, Japan has an Individualistic vs. Collectivist score of 62, meaning that Japan is an individualistic country; in terms of granularity, Japan is more individualistic than the United States, which has an Individualistic vs. Collectivist score of 60. However, each LLM we tested along with each approach we tested perceived the United States as predominantly individualistic and Japan as predominantly collectivist, with the largest discrepancy being within the personas approach for Command-R, where 72.40% of responses for the American persona were individualistic and only 19.60% of answers for the Japanese persona where individualistic. This may be because much of English language data represents Japan as a collectivist country (Scroope, 2021) and the United States as an individualistic country (Rosenbaum et al., 2018), leading to stereotypical representations of each country rather than true representations according to their Hofstede cultural dimensions. These findings hold for other individualistic countries often perceived as collectivist, such as South Korea (Yuh, 2016).
@@ Appendix E Full Data and Visualizations
Figure 3. GPT4o adhering well to individualism vs. collectivist value for high-resource languages
@@
Table 4. Language and Hofstede Cultural Dimensions Metadata
@@
Table 5. Graphs showing value differentiation across all models, approaches, and values. Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values.
@@
Table 6. Graphs showing value differentiation across all models, approaches, and values (continuation). Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
@@
Table 7. Graphs showing value differentiation across all models, approaches, and values (continuation). Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
@@
Table 8. Graphs showing value differentiation across all models, approaches, and values (continuation). Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
@@
Table 9. Graphs showing value differentiation across all models, approaches, and values. Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
@@
Table 10. Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs.
@@
Table 11. Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
@@
Table 12. Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
@@
Table 13. Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
@@
Table 14. Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
@@
Table 15. Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs.